# Physically Valid Biomolecular Interaction Modeling with Gauss-Seidel Projection

### v1

@@ 4.3 Derivatives via Implicit Differentiation
Figure 3: Importance of differentiable projection and finetuning. A post-hoc projection without finetuning applied to a 22-step sampling, while ensuring physical validity, fails to recover the $\alpha$-helical secondary structure (left). Integrating the projection as a differentiable layer and finetuning the diffusion module restores the helix and improves overall structural accuracy (right).
@@
Backpropagating through the Gauss-Seidel projection is non-trivial because it is implemented with an iterative solver. A naive approach, such as unrolling the forward iterations for automatic differentiation, is prohibitively memory-intensive and thus impractical for end-to-end training. We resort to implicit differentiation, a technique from sensitivity analysis (Burczyński et al., 1997). Specifically, we differentiate the first-order optimality condition in Eq. 2 with respect to $\hat{\mathbf{x}}$:
@@
Analogously to the forward pass, the backward pass requires solving an additional linear system. We use conjugate gradients (CG) to solve Eq. 8. whose convergence is guaranteed near feasibility (see Appendix F for details). For runtime efficiency, we implement the CG solver on the GPU. Algorithm 2 outlines the backward process of the projection module.
Table 1: Quantitative results on 3 of 6 datasets. Our method uses only 2 denoising steps yet always guarantees physical validity and achieves competitive structural accuracy compared to baselines. Dark green cells denote the best results among the few-step methods, while light green cells denote the best results among the 200-step methods. Results for the remaining datasets are in Appendix A.
[…]
CASP15
@@
Test
Boltz-1 200 0.80±0.01 0.54±0.01 0.48±0.03 0.69±0.06 0.58±0.06 0.56±0.06 0.44±0.02
Protenix 200 0.79±0.01 0.53±0.01 0.38±0.02 0.68±0.06 0.47±0.05 0.42±0.05 0.46±0.02
Boltz-1-Steering 200 0.80±0.01 0.53±0.01 0.51±0.03 0.67±0.06 0.58±0.05 0.55±0.06 0.99±0.00
[…]
Protenix-Mini 5 0.72±0.02 0.41±0.01 0.35±0.02 0.55±0.06 0.43±0.05 0.36±0.05 0.38±0.02
[…]
PoseBusters
Boltz-1 200 0.92±0.01 0.85±0.01 0.67±0.01 0.60±0.10 0.69±0.04 0.62±0.05 0.21±0.01
Protenix 200 0.92±0.01 0.87±0.00 0.72±0.01 0.63±0.11 0.74±0.04 0.67±0.05 0.37±0.03
[…]
5 Evaluation
In this section, we evaluate our method’s structural accuracy and physical validity through a series of experiments. We conduct quantitative results on six protein-complex benchmarks using standard evaluation metrics (Sec. 5.2). We also provide qualitative visualizations that compare protein structures and demonstrate the physical validity (Sec. 5.3). Furthermore, we analyze the convergence speed and wall-clock runtime of our approach (Sec. 5.4). Finally, we conduct an ablation study to isolate the impact of our Gauss-Seidel projection (Sec. 5.5).
5.1 Implementation Details
With the differentiable Gauss-Seidel projection module in place, we finetune a pre-trained all-atom diffusion network specifically for $2$-step sampling. During finetuning, we retain the original noise parameterization and denoising objective. At inference time, we use the inference-time sampling schedule of Protenix-Mini (Team et al., 2025). A Gauss-Seidel projection is applied to the model’s output to guarantee a physically valid final structure. Appendix C provides pseudo-code and complete algorithms for both training and inference with our projection module. We implement the differentiable Gauss-Seidel projection on the GPU using the WARP library (Macklin et al., 2024). In the forward pass, constraints are partitioned into batches with no shared atoms, allowing them to be processed in parallel. Coordinate updates are accumulated with atomic operations. We use T=20T{=}20 sweeps. The backward pass solves the adjoint linear system with a conjugate gradient solver fully on GPU, parallelizing per-constraint operations with atomic accumulation. We set the tolerance to $\varepsilon{=}10^{-4}$ and the maximal number of iterations to 1000. The projection module integrates with PyTorch as a custom autograd layer, avoiding the overhead costs of host-device copies and context switching. For training, we finetune Boltz-1 (Wohlwend et al., 2024) from open-source pretrained weights on the RCSB dataset (Burley et al., 2022) ($\sim\!180{,}000$ structures) using 8 NVIDIA H100 GPUs over two weeks. All evaluations are conducted on NVIDIA A6000 GPUs. For baselines and evaluation protocol, see Appendix H for details.
5.2 Quantitative Results
Table 1 and 2 (in Appendix A) summarize the comparison across six protein-complex benchmarks. Among all methods, ours uses the fewest denoising steps ($2$ steps). It achieves state-of-the-art accuracy in the few-step regime: Compared to Protenix-Mini ($5$ steps), our model improves complex- and interface-level LDDT across all benchmarks. Compared to $200$-step baselines, the accuracy of our method is competitive; on CASP15 and PoseBusters, ours outperforms Boltz-1 and Protenix on multiple metrics while approaching Boltz-2, which requires $100$ times more sampling. The largest gains appear in Lig-Prot LDDT, where our method leads on CASP15 and TEST and remains close to the best on PoseBusters, consistent with the projection module’s focus on atom-level geometry.
5.3 Qualitative Results
Fig. 4 and 6 (in Appendix B) present qualitative comparisons, showing that our approach consistently yields physically valid structures while preserving high accuracy. For PDB 7Y9A (with an N-acetyl-D-glucosamine ligand), which is used to assess ligand–protein interface quality, all baselines produce severe ligand–protein atom clashes, whereas our method eliminates these violations. For PDB 7XYO ($557$ residues with extensive secondary structure), which is used to evaluate large, structured assemblies, Boltz-1, Boltz-1-Steering, and Protenix show inter-chain collisions. Protenix-Mini avoids clashes but introduces large distortion, particularly in the second coiled segment. In contrast, our method produces clash-free structures while maintaining structural accuracy. Additional examples are provided in Appendix B.
Figure 4: Qualitative comparison with baseline methods. The red color highlights physically invalid predictions, such as atomic clashes. Our approach consistently guarantees physical validity.
5.4 Analysis
Wall-clock Runtime Comparison. To assess practical speed, we measured the wall-clock inference time (including diffusion and physical validity enforcement) on CASP15. The results, binned by atom count, are shown in Fig. 5 (right). Our $2$-step predictor with Gauss-Seidel projection is consistently the fastest. It achieves a median speedup of ${\sim}9.4\times$ over the 200‑step Boltz‑1 baseline and ${\sim}9.5\times$ over Protenix‑Mini. Compared with Boltz‑1-Steering, the speedup is $23{-}46\times$. The gains stem from the reduction in denoising steps and our GPU implementation of the projection module.
5.5 Ablation Study
Complex [-0.15ex]LDDT Physical [-0.15ex]Validity
200-step $0.6204$ $0.6209$
(i) + 2-step (no guidance, no finetune) $0.5814$ $0.5846$
(ii)    + GD at inference, no finetune $0.5778$ $0.7050$
(iii)    + GS at inference, no finetune $0.5768$ $\mathbf{1.00}$
(iv)     + GS at inference, finetune $\mathbf{0.6239}$ $\mathbf{1.00}$
Figure 5: Convergence and runtime. Left: Potential energy vs. iteration on PDB 8X51 under four noise levels ($\sigma\!\in\!\{160,120,80,40\}$). Gauss-Seidel projection (orange) converges to tolerance within 2020 iterations; gradient descent guidance (blue) oscillates and decays slowly. Right: Wall-clock inference time by atom-count bin on CASP15. Ours is ${\sim}10\times$ faster than baselines while maintaining validity. Runtime includes diffusion and validity module; measured on the same hardware.
6 Conclusion
Reproducibility Statement.
To ensure the reproducibility of our work, we provide detailed experimental settings in Sec. 5.1, the pseudocode for the Forward Solver and Backward Pass of the Gauss-Seidel projection module in Sec. 4.2 and Sec. 4.3, the pseudocode for training and inference algorithms in Appendix C, the details on the baselines and the evaluation protocol in Appendix H, and the complete formulations of the physical validity constraints in Appendix G.
Appendix A Additional Quantitative Results
Table 2 shows the quantitative comparison on AF3-AB, dsDNA, and RNA-Protein datasets. Our method uses only two denoising steps yet always guarantees physical validity and achieves competitive structural accuracy compared to baselines.
Table 2: Quantitative results on the remaining datasets. Dark green cells denote the best results among the few-step methods. Light green cells denote the best results among the 200-step methods.
Method Metric # Denoise Steps Complex [-0.15ex]LDDT DNA-Prot [-0.15ex]LDDT RNA-Prot [-0.15ex]LDDT Physical [-0.15ex]Validity
AF3-AB
Boltz-1 200 0.80±0.02 0.00±0.00
Protenix 200 0.79±0.03 0.00±0.00
Boltz-Steering 200 0.79±0.03 0.90±0.02
[…]
Protenix-Mini 5 0.75±0.03 0.00±0.00
[…]
dsDNA
Boltz-1 200 0.85±0.04 0.76±0.02 0.03±0.01
Protenix 200 0.86±0.04 0.78±0.02 0.06±0.01
[…]
RNA-Protein
Boltz-1 200 0.74±0.05 0.30±0.02 0.00±0.00
@@ Appendix A Additional Quantitative Results
Appendix B Additional Qualitative Results
@@
Fig. 6 shows more qualitative comparisons with baselines. We also visualize full protein structures predicted by our method and compare them with the ground truth from six datasets, shown in Fig. 7. The results demonstrate that our model is capable of accurately predicting physically valid structures.

### v2

@@ 4.3 Derivatives via Implicit Differentiation
Figure 3: Importance of differentiable projection and finetuning. A post-hoc projection without finetuning with a 22-step sampling, while ensuring physical validity, fails to recover the $\alpha$-helical secondary structure (left). Integrating the projection as a differentiable layer and finetuning diffusion module restores the helix and improves structural accuracy (right).
@@
Backpropagating through the Gauss-Seidel projection is non-trivial because it is implemented with an iterative solver. A naive approach, such as unrolling the forward iterations for automatic differentiation, is prohibitively memory-intensive and thus impractical for end-to-end training. We resort to implicit differentiation, a technique from sensitivity analysis (burczynski1997comparison). Specifically, we differentiate the first-order optimality condition in Eq. 2 with respect to $\hat{\mathbf{x}}$:
@@
Analogously to the forward pass, the backward pass requires solving an additional linear system. We use conjugate gradients (CG) to solve Eq. 8, whose convergence is guaranteed near feasibility (see Appendix K for details). For runtime efficiency, we implement the CG solver on the GPU. Algorithm 2 outlines the backward process of the projection module.
5 Evaluation
In this section, we evaluate our method’s structural accuracy and physical validity through a series of experiments. We report quantitative results on six protein-complex benchmarks using standard evaluation metrics (Sec. 5.2). We also provide qualitative visualizations that compare protein structures and demonstrate the physical validity (Sec. 5.3). Furthermore, we analyze the convergence speed and wall-clock runtime of our approach (Sec. 5.4). Finally, we conduct an ablation study to isolate the impact of our Gauss-Seidel projection (Sec. 5.5).
5.1 Implementation Details
With the differentiable Gauss-Seidel projection module in place, we finetune a pre-trained all-atom diffusion network specifically for $2$-step sampling. During finetuning, we retain the original noise parameterization and denoising objective. At inference time, we use the inference-time sampling schedule of Protenix-Mini (bytedance2025protenix). A Gauss-Seidel projection is applied to the model’s output to guarantee a physically valid final structure. Appendix H provides pseudo-code and complete algorithms for both training and inference with our projection module. We implement the differentiable Gauss-Seidel projection on the GPU using the WARP library (warp2022). In the forward pass, constraints are partitioned into batches with no shared atoms, allowing them to be processed in parallel. Coordinate updates are accumulated with atomic operations. We use T=20T{=}20 sweeps. The backward pass solves the adjoint linear system with a conjugate gradient solver fully on GPU, parallelizing per-constraint operations with atomic accumulation. We set the tolerance to $\varepsilon{=}10^{-4}$ and the maximal number of iterations to 1000. The projection module integrates with PyTorch as a custom autograd layer, avoiding the overhead costs of host-device copies and context switching. For training, we finetune Boltz-1 (wohlwend2024boltz1) from open-source pretrained weights on the RCSB dataset (burley2022rcsb) ($\sim\!180{,}000$ structures) using 8 NVIDIA H100 GPUs over two weeks. All evaluations are conducted on NVIDIA A6000 GPUs. For baselines and evaluation protocol, see Appendix M for details.
Table 1: Quantitative results on 6 datasets (Part I). Our method uses only 2 denoising steps yet always guarantees physical validity and achieves competitive structural accuracy compared to baselines. Dark green cells denote the best results among the few-step methods, while light green cells denote the best results among the 200-step methods. We also include the results of AlphaFold3 (AlphaFold3) for references.
@@ 5.1 Implementation Details
[…]
5.2 Quantitative Results
Table 1 and 2 summarize the comparison across six protein-complex benchmarks. Among all methods, ours uses the fewest denoising steps ($2$ steps). It achieves state-of-the-art accuracy in the few-step regime: Compared to Protenix-Mini ($5$ steps), our model improves complex- and interface-level LDDT across most benchmarks. Compared to $200$-step baselines, the accuracy of our method is competitive. On Test and PoseBusters, our performance is comparable to Boltz-1 and Protenix across multiple metrics while approaching Boltz-2, which requires $100$ times more sampling.
5.3 Qualitative Results
Fig. 4 and 9 (in Appendix A) present qualitative comparisons, showing that our approach consistently yields physically valid structures while preserving high accuracy. For PDB 7Y9A (with an N-acetyl-D-glucosamine ligand), which is used to assess ligand–protein interface quality, all baselines produce severe ligand–protein atom clashes, whereas our method eliminates these violations. For PDB 7XYO ($557$ residues with extensive secondary structure), which is used to evaluate large, structured assemblies, Boltz-1, Boltz-1-Steering, and Protenix show inter-chain collisions. Protenix-Mini avoids clashes but introduces large distortion, particularly in the second coiled segment. In contrast, our method produces clash-free structures while maintaining structural accuracy. Additional examples are provided in Appendix A.
5.4 Analysis
Figure 4: Qualitative comparison with baseline methods. The red color highlights physically invalid predictions, such as atomic clashes. Our approach consistently guarantees physical validity. The first example demonstrates a ligand-protein steric clash, where the ligand atoms physically penetrate the receptor surface. This corresponds to constraint 1 (Steric Clash) in Fig. 11. The second example demonstrates an inter-chain overlap, where atoms from Chain A and Chain B occupy the same space. This corresponds to constraint 6 (Overlapping Chains) in Fig. 11.
Wall-clock Runtime Comparison. To assess practical speed, we measured the wall-clock inference time (including diffusion and physical validity enforcement) on CASP15. The results, binned by atom count, are shown in Fig. 5 (right). Our $2$-step predictor with Gauss-Seidel projection is consistently the fastest. It achieves a median speedup of ${\sim}9.4\times$ over the 200‑step Boltz‑1 baseline and ${\sim}9.5\times$ over Protenix. Compared with Protenix-Mini with $5$ denoising steps, our method achieves ${\sim}2.3\times$ speedup. Compared with Boltz‑1-Steering, the speedup is $23{-}46\times$. The gains stem from the reduction in denoising steps and our GPU implementation of the projection module.
5.5 Ablation Study
Complex [-0.15ex]LDDT Physical [-0.15ex]Validity
200-step $0.6406$ $0.6545$
(i) + 2-step (no guidance, no finetune) $0.5920$ $0.6327$
(ii)    + GD at inference, no finetune $0.5942$ $0.7345$
(iii)    + GS at inference, no finetune $0.5889$ $\mathbf{1.00}$
(iv)     + GS at inference, finetune $\mathbf{0.6321}$ $\mathbf{1.00}$
Figure 5: Convergence and runtime. Left: Potential energy vs. iteration on PDB 8X51 under four noise levels ($\sigma\!\in\!\{160,120,80,40\}$). Gauss-Seidel projection (orange) converges to tolerance within 2020 iterations; gradient descent guidance (blue) oscillates and decays slowly. Right: Wall-clock inference time by atom-count bin on CASP15. Ours is ${\sim}10\times$ faster than 200-step baselines and ${\sim}2.3\times$ faster than Protenix-Mini, while maintaining validity. Runtime includes diffusion and validity module; measured on the same hardware.
Table 2: Quantitative results on 6 datasets (Part II). Dark green cells denote the best results among the few-step methods. Light green cells denote the best results among the 200-step methods.
Method Metric # Denoise Steps Complex [-0.15ex]LDDT Prot-Prot [-0.15ex]iLDDT DNA-Prot [-0.15ex]iLDDT RNA-Prot [-0.15ex]iLDDT TM-score Physical [-0.15ex]Validity
AF3-AB Boltz-1 200 0.80±0.02 0.12±0.05 0.61±0.06 0.00±0.00
Protenix 200 0.80±0.03 0.23±0.06 0.64±0.06 0.01±0.01
[…]
Protenix-Mini 5 0.77±0.03 0.11±0.04 0.59±0.06 0.00±0.00
[…]
dsDNA Boltz-1 200 0.85±0.04 0.76±0.07 0.89±0.06 0.03±0.03
Protenix 200 0.86±0.04 0.78±0.08 0.89±0.06 0.04±0.03
[…]
RNA-Protein Boltz-1 200 0.74±0.04 0.30±0.07 0.85±0.11 0.00±0.00
Protenix 200 0.75±0.05 0.34±0.09 0.82±0.12 0.02±0.02
[…]
Protenix-Mini 5 0.72±0.05 0.36±0.10 0.79±0.13 0.05±0.05
[…]
6 Conclusion
Reproducibility Statement.
To ensure the reproducibility of our work, we provide detailed experimental settings in Sec. 5.1, the pseudocode for the Forward Solver and Backward Pass of the Gauss-Seidel projection module in Sec. 4.2 and Sec. 4.3, the pseudocode for training and inference algorithms in Appendix H, the details on the baselines and the evaluation protocol in Appendix M, and the complete formulations of the physical validity constraints in Appendix L.
Appendix A Additional Qualitative Results
@@
Fig. 9 shows more qualitative comparisons with baselines. We also visualize full protein structures predicted by our method and compare them with the ground truth from six datasets, shown in Fig. 10. The results demonstrate that our model accurately predicts physically valid structures.
Appendix B Uniqueness of the Gauss-Seidel Solution
The Gauss-Seidel scheme guarantees uniqueness for linearized subproblems. Theoretically, while ordering and initialization govern the convergence speed of the Gauss-Seidel scheme, they do not influence the solution’s uniqueness (golub2013matrix).
We investigated this empirically on the PoseBusters dataset by performing five independent projection runs for each denoised structure. The computed pairwise RMSD among outputs was 0.0000 $\pm$ 0.0000 Å. This confirms that, in practice, the projection module behaves as a stable mapping, reliably converging to a numerically identical solution.
Appendix C Practical choice of $\alpha$ and its sensitivity of convergence
To assess how the parameter $\alpha$ influences the convergence of the Gauss-Seidel projection, we conducted an ablation study on the protein PDB 7WUY, evaluating both convergence behavior and predicted accuracy across six values ($\alpha{\in}\{10^{-2},\dots,10^{-7}\}$). As illustrated in Fig. 6, we monitored the potential energy after each iteration and observed two distinct behaviors: For large $\alpha$ ($10^{-2},10^{-3},10^{-4}$), it fails to converge to a physically valid configuration. While the potential energy decreases initially, it exhibits persistent oscillations rather than converging to zero. This indicates that the constraint penalties are not enforced strongly enough to resolve physical violations. For small $\alpha$ ($\leq 10^{-5}$), at $\alpha{=}10^{-5}$, the energy oscillates transiently before eventually converging to zero, confirming that all constraints are resolved. For smaller values ($\alpha{=}10^{-6},10^{-7}$), the projection consistently and rapidly converges to a valid state (zero potential energy), demonstrating fast and stable constraint satisfaction. These results confirm that selecting $\alpha{\leq}10^{-6}$ is critical for ensuring the projection reliably maps to a physically valid configuration.
Figure 6: Convergence and computational cost analysis of the Gauss-Seidel projection module. Left: Potential energy plotted over iterations for $\alpha$ values ranging from $10^{-2}$ to $10^{-7}$. Large values ($10^{-5}$) result in persistent oscillations without full convergence. An intermediate value ($\alpha{=}10^{-5}$) exhibits transient oscillations before reaching zero potential, while smaller values ($10^{-6},10^{-7}$) ensure rapid and stable convergence, indicating robust constraint satisfaction. Right: Computational benchmarks showing forward/backward time costs and additional memory overhead across varying token counts (256, 384, 512, and 640).
Figure 7: Comparison of Boltz-1, Boltz-2, and our method across 2-, 5-, 10-, and 200-step budgets. Our method consistently outperforms the Boltz-1 baseline while maintaining strong physical validity. Although Boltz-2 achieves higher Complex-LDDT scores, integrating our module into the Boltz-2 preserves this accuracy while significantly boosting physical validity at low step budgets.
Figure 8: Visualization of structural corrections by Guass-Seidel projection. The projection effectively resolves steric violations and restores physically valid geometries. We showcase four representative cases from the PoseBusters dataset. PDB 7P4C: A distorted $\alpha$-helix clashing with a neighboring strand is corrected. PDB 7PL1: A self-entangled coil is disentangled, resolving clashes. PDB 7ELT: Ligand-sidechain interactions and invalid internal ligand geometries are resolved. PDB 8AEM: Ligand-related collisions are removed. For each case, RMSD values were computed to quantify the structural refinement.
@@ Appendix C Practical choice of $\alpha$ and its sensitivity of convergence
Appendix D Comparisons with different denoising step budgets
We conducted experiments comparing Boltz-1, Boltz-2, and our method (built on Boltz-1) across denoising step budgets of 2, 5, 10, and 200 on the CASP15 dataset. As shown in Fig. 7, our method demonstrates consistent improvement over the Boltz-1 baseline, outperforming it in terms of both structural accuracy (Complex-LDDT) and physical validity across all step budgets.
We acknowledge that our current implementation does not surpass the performance of Boltz-2. We attribute this to the underlying base model: our method is fine-tuned using the Boltz-1 architecture, weights, and dataset. In contrast, Boltz-2 benefits from significant architectural advancements (e.g., the affinity prediction module) and a much larger training dataset particularly focused on protein complex interactions. This advantage is evident in the interface metrics where Boltz-2 dominates.
However, our Gauss-Seidel projection module is theoretically agnostic to the base model. To validate this potential without the computational cost of full fine-tuning, we applied our module as a post-processing step on top of Boltz-2. Results in Fig. 7 show that this hybrid approach effectively ensures 100% physical validity: it resolves invalid outputs observed in Boltz-2 at low step counts (<<10) while maintaining the high structural accuracy of the original model. We thus treat our method as a general-purpose physical validity enforcement module that can be flexibly integrated into various diffusion-based structure prediction frameworks.
Appendix E Computational Cost of the Gauss-Seidel Projection Module
We systematically evaluated the computational overhead of incorporating the Gauss-Seidel projection module during both training and inference. To provide a canonical measure of input complexity, we benchmarked the module on the CASP15 dataset, binning samples by token count (256, 384, 512, and 640). We reported the forward and backward wall-clock runtimes and additional memory usage on an NVIDIA A100 GPU.
As illustrated in Fig. 6, both forward and backward passes are completed within a matter of seconds, and the additional memory overhead remains below 3.5 GB, even for the largest token bins. This high efficiency is achieved through our specialized implementation, which leverages custom CUDA kernels optimized for both the forward and backward processes.
Table 3: L-RMSD < 2 Å on the PoseBusters dataset computed by PXMeter (PXMeter). Dark green cells denote the best results among the few-step methods. Light green cells denote the best results among the 200-step methods.
Dataset Method AlphaFold3 Boltz-1 Protenix Boltz-1-Steering Boltz-2 Protenix-Mini [-0.15ex]10-step Protenix-Mini [-0.15ex]5-step Ours
[…]
Appendix F Clarification on Docking-related Metric Computation
In our primary evaluation, we calculated docking-related metrics (DockQ >> 0.23, and L-RMSD << 2Å) using OpenStructure (biasini2013openstructure), strictly adhering to the protocol defined in the original Boltz-1 paper (wohlwend2024boltz1). We attribute the discrepancies between our reported values and those in the Protenix technical report (bytedance2025protenix) to their use of a different evaluation tool, PXMeter (PXMeter). To investigate this, we performed an evaluation of the L-RMSD << 2Å metric using PXMeter on the PoseBusters dataset for all methods (detailed in Table 3). This re-evaluation confirmed two key points: 1) The re-calculated performance of the Protenix baseline aligns closely with the figures in their original report, validating the source of the numerical difference. 2) While the absolute values depending on the tool used, the relative ranking among the baselines remains unchanged. These results indicate that our comparative analysis is informative.
Appendix G Analysis of Structural Perturbations Introduced by Gauss-Seide Projection
We quantified the structural changes introduced by Gauss-Seidel projection by computing the RMSD between pre- and post-projection structures on the PoseBusters dataset. The observed average RMSD was 0.1865 $\pm$ 0.0219 Å, confirming that the projection step introduces atomic position changes sufficient to enforce physical validity. Fig. 8 visualizes four representative cases where noticeable improvements occurred. For PDB 7P4C, the projection restored a distorted $\alpha$-helix that was originally clashing with a neighboring strand. In PDB 7PL1, a self-entanglement within a coil region was successfully disentangled. Similarly, for PDB 7ELT and PDB 8AEM, the module resolved ligand-sidechain and intra-ligand steric clashes. These examples demonstrate that the Gauss-Seidel projection effectively enforces physical validity while preserving the structural prediction of the generative model.